I am unsure as to how to begin to solve this word problem. The OfficeJet printer can copy Maria's dissertation in 8 minutes. The LaserJet printer can copy the same document in 12 minutes. If the two machines worked together, how long would they take to copy the dissertation? Thank you for any help you can provide on this one.
Answer by stanbon(57412)

The OfficeJet printer can copy Maria's dissertation in 8 minutes. The LaserJet printer can copy the same document in 12 minutes. If the two machines worked together, how long would they take to copy the dissertation?
-----
Office Jet time = 8 min/job ; rate = 1/8 job/min
-----
Laser Jet time = 12 min/job ; rate = 1/12 job/min
-----
Together time = x min/job ; rate = 1/x job/min
---------------------------------
Equation:
rate + rate = together rate
1/8 + 1/12 = 1/x
----
Multiply thru by 24x to get:
3x + 2x = 24
5x = 24
x = 4.8 minutes/job (time to do the job together)
=====================================================
Cheers,
Stan H.
=================